Solve using completing the square,
3x^2-6x=27

Answer by stanbon(48569)

Solve using completing the square,
3x^2-6x=27
-----------
1st: Simplify:
x^2 - 2x = 9
x^2-2x+1 = 10
(x-1)^2 = 10
x-1 = sqrt(10) or (x-1) = -sqrt(10)
x = 1+sqrt(10) or x = 1-sqrt(10)
===================================
Cheers,
Stan H.